Minutes of the Tracker Meeting

July 8, 2004

 

Action Items:

  1. Arthur: investigate next week the issues with the cable bending tool
  2. Dave: include a load proof test of the inner shipping container
  3. Jeff: get an improved battery life for the environmental monitor used in tower shipping
  4. Martin: send the cable mounting bracket designs to Chris
  5. Dave and Jeff: keep some of the M2.5 screws and M5 hex nuts at SLAC for evaluation
  6. Travelers in Pisa: bring back shipping containers for re-use
  7. Jack?: see if a TQCM can be borrowed from NASA
  8. Sandro: sign the sidewall drawings
  9. Martin: sign the G&A procedure

 

Status of Alignment Activitiesa: The CMM was reconfigured to increase its travel range.Tomorrow morning they will be able to try out the alignment procedure.

 

Sandro noted that they did some tests with the cones and found that they could achieve a setting error of +-10 microns.

 

Status of SLAC deliveries for next week:

- Reworked MCMs: 45 returned from Zentek yesterday, with the new 75 ohm resistors.They are going through testing again and at least 36 will go to Pisa with Robert on Monday.

- Cable tester: SLAC has 2 units, so Robert will bring one to Italy.

- Lifting fixture: it is in the SLAC shop and on track for a Friday delivery.Dave is working with the riggers for a proof load certification on Saturday.John Ku and Mike Foss are involved.

- Inner shipping container: it is at an outside shop and scheduled for delivery tomorrow.Martin noted that this also needs a load test, since it is also used to lift the tower.It needs a cable holding plate for assembly.Probably it will ship to Pisa next week, not be hand carried.

- Cable holding plate: there was a problem with purchasing keeping the vendor on hold.The machine shop is now working on the first article, which will be available Saturday.They had to skip the alodyne finish to make that shedule.CMM of this is scheduled for Saturday.

- Interface hardware: 18 sets of studs, cones, and hex bushings are at SLAC.Nytemp is on the studs.Arthur validated the process and checked the fit and running torque.The studs and bushings are currently out for Dycronite coating and will be back tomorrow at noon.The M5 hex nuts are complete and will be delivered tomorrow.

- Cable bending tool: the radii don't look correct.There is an NCR, and this needs to be evaluated and corrected next week.

- Outer shipping container: Ku and Foss are working on it with Jeff.Mike Opie will look at the current load design.They are coordinating with Tom Borden, who supervised testing of the previous container.They should email all the info to INFN and Martin Nordby before ordering.Sandro needs the dimensions immediately in order to size the truck needed to cart the tower around Italy.Jeff said it will be ength 37.75" x width 37.75" x height 48.00", all outside diameter.This keeps the same foot print as the previous container but lowers the height by 11".It was decided that there should be no castors.Instead, it should be bolted to a plastic pallet.Jeff noted that the height leaves 6 inches clear inside the container top and bottom.

It was noted that we need a longer battery life for the environmental recorder, which only worked for 4 days last time.

- Corner cable mounting brackets: the design is just being completed by Foss.A PR was submitted for quotes.SLAC will pay an expediting fee for the first units, but there is no delivery date yet.Martin should send the design info to Chris Fransen.

- heat straps: these are in fab at SLAC with delivery Friday afternoon; CMM is scheduled for Saturday.

- torque wrenches: Dave procured a full complement of preset torque drivers; 2 of each.Martin said he still needs to review the torque specs. Arthur said that we also need another torque meter.

- Torx screwdrivers and bits were also obtained.

- M2.5 100-degree sidewall fasteners: the screws are due tomorrow.SLAC should keep some for inspections; SLAC should also keep some of the hex nuts.

- The new vibe fixture will come in later this month; we must use the existing fixture for Tower A

- Dave requested that all travelers bring shipping cases back from Pisa!

- TQCM: GSFC and Jack insist that it is needed.Nanda is checking more in Italy.Can one be obtained from NASA?

- Reflector balls and nests: the PR went in yesterday.

 

Flight cable status: Darren reported that 16 cables went to INFN.Parlex could not produce coupons for one cable.Robert said that the 2nd set really should go to the electronics group at SLAC and would be non-flight at that point. The plan is to keep both sets in Italy for Tower-A, to make sure that spares are available, and then ship the ones not used to Gunther.The 3rd and forth sets were due July 16 but are delayed a couple of weeks due to various issues.The coupons will go to Richard Gobin and then to NASA for testing.

 

Mid Tray Panel Fabrication: Sandro reported on the coupon tests at Plyform.The bead blasting with 0.5um texture increased adhesion by a factor of 2.7 with respect to the old hand sanding process.Monday Sandro will go to Bologna, where they will prepare tungsten for 1 tray with 0.5um roughness plus a few other samples with other higher roughness specs.Plyform will then make 1 more tray and will do more coupon tests.Sandro will also on Monday book the company for the remaining plates.The company will need to make a simple tool for the foils to go into an automatic machine.When they are finished they will pack them with nitrogen.Sandro will not know the schedule until Monday.

- MRB for the Kapton-Tungsten bonding issue: a plan was hatched to hold it in Italy next week.Ben Rodini is available Thursday morning EDT.

 

Bottom and Top Tray Panel Fabrication: Riccardo reported that the Tower A tray was assembled and measured last week.He has been reviewing the results with Martin, and they are improving the metrology procedure.There wonít be time this week, however, to remeasure the tray, so Riccardo will do it on Monday.From the existing data the flexure positions appear not to be within the drawing tolerances.After the remeasure Martin will figure out whether this really is an issue.Riccardo also reported that Plyform finished the top tray closeout machining.Riccardo will go tomorrow to G&A to pick up the tooling for top tray assembly.Plyform will assemble it next week.

 

G&A Tray Integration: Sandro reported that today G&A should have assembled MCMs onto 2 trays with a redlined procedure.The modification is to use a washer and nut on each of the 3 grounding pins.This is needed in order to make good electrical contact to the ground pad on the front side of the MCM.Tomorrow they will do the assembly of ladders, followed by wire bonding on the weekend.Luca and Luca will go to G&A on Monday to install the test setup and execute the test.Tuesday G&A will do the encapsulation, which then requires a long cure.The following Monday Sandro plans to test the trays with encapsulation.They will do a T/V test at G&A to 50C for a few hours, with the tray in its shipping container.

- drawing and procedure release: The drawings have been signed off but now need to be modified with the new nuts and washers.The procedure document lacks just Martinís signature, and also will need modification for the nuts and washers.The tray test document is in release.

 

Sidewall Fabrication: a detailed schedule will be obtained tomorrow from Plyform.Martin noted that we still need Sandro's signature on the drawings.

 

EM tower shipping to I&T and the Tower-A shipping container: we want to leave the existing shipping container in Italy, so the EM tower will be put into its inner container (which needs the vibe fixture from the tower mockup) and will be crated with lots of foam padding.

 

The mockup tower may be needed in Pisa through next week.We will review this after tomorrow's experience.

 

Arthur noted that he saw galvanic corrosion between shims and the aluminum grid simulator.He said that we really need coating of the shims as well as the grid.

 

Sandro noted that they did do some measurements of the EM tower before and after vibration.They saw only a 50 micron maximum deviation.Of course, this was not done with the standard metrology program being installed this week.